OpenStudy (anonymous):

1. find the distance between points at (-1, 6) and (5,-2) 2. find the distance between points at (-5,2) and (7,-3)

OpenStudy (anonymous):

\[\sqrt{(y2-y1)^{2} +(x2-x1)^{2}}\]

OpenStudy (anonymous):

\[\sqrt{(-2-6)^{2}+(5-(-1)^{2}}\]

OpenStudy (jdoe0001):

\(\bf \textit{distance between 2 points}\\ \quad \\ \begin{array}{lllll} &x_1&y_1&x_2&y_2\\ % (a,b) &({\color{red}{ -1}}\quad ,&{\color{blue}{ 6}})\quad % (c,d) &({\color{red}{ 5}}\quad ,&{\color{blue}{ -2}})\\ &({\color{red}{ -5}}\quad ,&{\color{blue}{ 2}})\quad % (c,d) &({\color{red}{ 7}}\quad ,&{\color{blue}{ -3}}) \end{array}\qquad % distance value d = \sqrt{({\color{red}{ x_2}}-{\color{red}{ x_1}})^2 + ({\color{blue}{ y_2}}-{\color{blue}{ y_1}})^2}\)

OpenStudy (anonymous):

\[\sqrt{(-8)^{2}+(6)^{2}}\]

OpenStudy (anonymous):

\[\sqrt{(64)+(36)}\]

OpenStudy (anonymous):

\[\sqrt{100}\]

OpenStudy (anonymous):

10 ans

OpenStudy (anonymous):

now you can dothe number 2 exactly same way and get an answer

OpenStudy (anonymous):

ok give me a min

OpenStudy (anonymous):

if you get 13 as your answer, then you are correct. bye

OpenStudy (anonymous):

please close the question or maybe fan me too

OpenStudy (anonymous):

thank you